# Scaling Behaviors of LLM Reinforcement Learning Post-Training: An Empirical Study in Mathematical Reasoning

### v1

@@ Abstract
While scaling laws for large language models (LLMs) during pre-training have been extensively studied, their behavior under reinforcement learning (RL) post-training remains largely unexplored. This paper presents a systematic empirical investigation of scaling behaviors in RL-based post-training, with a particular focus on mathematical reasoning. Based on 54 experiments across diverse model sizes and training settings, we characterize how model scale, data volume, and computational budget interact to shape performance. Our analysis leads to four key findings: ❶ Under a fixed computational budget, larger models trained for fewer steps consistently outperform smaller models trained for more steps. ❷ Given a fixed amount of training data, larger models achieve superior sample efficiency, yielding lower loss. ❸ In data-constrained regimes, repeated reuse of high-quality data proves highly effective, as final performance is primarily governed by the total number of optimization steps rather than the uniqueness of samples. ❹ These scaling behaviors are robust across both base and instruction-tuned models, which share similar learning dynamics (e.g., larger models show faster convergence) even while differing in absolute accuracy. Collectively, these results provide a principled foundation and practical guidelines for efficiently scaling the reasoning capabilities of LLMs through RL post-training.
@@ 1 Introduction
The rapid progress of large language models (LLMs) has made elucidating their scaling laws a matter of central importance. These laws, which capture the intricate relationships between model architecture, parameter size, computational cost, data availability, and downstream performance (Kaplan et al., 2020; Hoffmann et al., 2022), are invaluable not only because they illuminate the latent factors governing learning dynamics, but also because they provide actionable guidance on how to distribute scarce computational resources most effectively (Li et al., 2025a). While extensive efforts have clarified scaling behavior, the scaling behavior of reinforcement learning (RL) post-training for LLM reasoning remains underexplored.
@@
During pretraining, Kaplan et al. (2020) show that cross-entropy loss follows smooth power-law scaling in model size, dataset size, and training compute, implying that larger models trained for fewer steps are compute-optimal. Hoffmann et al. (2022) refine this by showing that, under fixed compute, scaling parameters and tokens proportionally is optimal, since many large models are undertrained. Extending to neural-based RL, Hilton et al. (2023) empirically demonstrates that the intrinsic performance of convolutional neural networks (CNNs) optimized via reinforcement learning also scales power-law with model capacity and environment interaction.
Recently, RL has become the predominant post-training strategy for enhancing the reasoning abilities of LLMs, particularly in mathematics, a domain that demands long-horizon, compositional reasoning (Ferrag et al., 2025; DeepSeek-AI, 2025; Kimi Team, 2025; Ahn et al., 2024). Given the recent works applying RL for LLM reasoning, it is crucial to understand how to scale RL training effectively for reasoning tasks. In this work, we aim to fill this research gap with an extensive empirical study on models of varying scales. Defining the normalized test error rate as Test Loss ($L$), we formalize the resource allocation challenge as a series of constrained optimization problems:
  1. 1.
    The Compute-Constrained Scenario: Given a fixed computational budget $C$, we seek the model size $N$ (and corresponding data allocation $D$) that minimizes the final loss:
  2. 2.
    The Data-Constrained Scenario: For a fixed amount of unique training data $D$, we aim to determine the model size $N$ that yields the lowest test loss under this data limitation:
  3. 3.
    The Data Reuse Scenario: With a fixed number of optimization steps $S$, we investigate the impact of the data reuse factor $\tau$—the number of times each unique example is revisited—on the final loss. In this setting, the total number of processed samples is constrained as $D_{\text{unique}}\times\tau=S_{\text{const}}$. Formally:
    $\underset{\tau,~C}{\operatorname*{arg\,min}}\;L(\tau,C)\quad\text{s.t.}\quad D_{\text{unique}}\times\tau=S_{\text{const}},$ (3)
@@
To empirically answer these questions, we fine-tune 54 LLMs with reinforcement learning on 50k mathematics problems, based on the Qwen2.5 model family (Qwen et al., 2025). Figure 1 shows that, within the 0.5-14B range, and tested in an environment close to the training set distribution, the loss reduction brought by RL follows an approximately log-linear trend with compute and data size. Importantly, larger models not only have better initial performance but also generally have more efficient computation and data utilization during the optimization process.
(a) Relation between test loss and training compute for base models with fitted curve.
(b) Relation between test loss and data size for base models with fitted curve.
Figure 1: Key finding of this study: In both cases, larger models consistently outperform smaller models at all points in training.
@@
We further analyze the data-constrained regime, where we demonstrate that data reuse is a highly effective strategy. We validate the generality of our findings through extensive ablation studies on both base and instruct model series. Besides, we also study the impact of the key hyperparameter rollout number in the GRPO algorithm (Shao et al., 2024). These investigations establish fundamental scaling relationships for RL post-training, providing a quantitative foundation and practical guidelines for resource-efficient model refinement.
@@
  • In our experiment scale, larger models, starting with stronger initial performance, consistently achieve better compute and data efficiency in RL post-training for mathematical reasoning.
@@
  • In math reasoning, scaling behaviors are robust across both base and instruction models and GRPO hyperparameters, providing actionable guidance for resource-efficient RL post-training.
@@ 2 Experimental Setup
We use the Qwen2.5 model family (0.5B, 1.5B, 3B, 7B, and 14B parameters) (Qwen et al., 2025), which shares the same architecture, so that parameter count is the only variable in our scaling analysis. All experiments are run with the VeRL framework (Sheng et al., 2024), a large-scale RL platform for LLMs ensuring consistency and reproducibility.
@@
$\mathcal{L}_{\text{GRPO}}=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\lvert o_{i}\rvert}\sum_{t=1}^{\lvert o_{i}\rvert}\Bigl\{\min\Bigl[\rho(\theta)\,\hat{A}_{i,t},\;\operatorname{clip}\!\Bigl(\rho(\theta),1-\varepsilon,\,1+\varepsilon\Bigr)\hat{A}_{i,t}\Bigr]-\beta\,\mathrm{D}_{\mathrm{KL}}\Bigr\}.$ (4)
@@
Dataset settings.
The training data is the mathematics subset of the guru-RL-92k dataset from the Reasoning360 project (Cheng et al., 2025), which is carefully curated through deduplication and difficulty filtering. We further sort the problems by increasing difficulty (decreasing pass rate) to enable curriculum learning. The evaluation data consists of two parts. To derive scaling laws, we use a held-out set of 500 in-domain math problems sampled from the training distribution. To assess generalization, we evaluate on a broader benchmark suite spanning mathematics (AIME2024 (Patel et al., 2024), AMC2023 (KnovelEng, 2025), GSM8K (Cobbe et al., 2021), MATH500 (Lightman et al., 2023)), code (HumanEval (Chen et al., 2021)), logic (Zebra Puzzle(Lin, 2024)), and science (SuperGPQA (Team et al., 2025)). More details about dataset settings can be found in Appendix A.1.
Prompt Setting.
To ensure stable behavior during RL training and evaluation, we use structured prompts tailored to each domain. For example, all mathematics problems are prepended with the Chain-of-Thought prompt (Wei et al., 2023): “You are a knowledgeable math assistant. Answer the following questions and think step by step”. More prompt templates for all related domains could be found in Appendix A.3.
@@
Our primary evaluation metric is the test loss ($L$), a proxy for reward-based performance in the RL setting. Unlike conventional RL, where agents interact with an environment and sample rewards freely, LLM post-training mainly focuses on the fixed data set and a verifiable reward rule. We hold out part of the training corpus and compute the normalized error rate. Formally, $L=1-(R/R_{\text{max}})$, where $R$ is the number of correct solutions and $R_{\text{max}}$ the total. Obviously, the goal of maximizing reward in RL training is equivalent to minimizing $L$.
@@ 3 Empirical Results and Scaling Laws
This section presents a comprehensive empirical investigation into the scaling behavior of RL for post-training LLMs. Our experiments are designed to address the core questions posed in Eq. 1, Eq. 2, and Eq. 3. We first examine scaling behaviors under compute and data constraints, then analyze independent scaling dimensions, data reuse strategies, and finally evaluate generalization performance together with an ablation study on the GRPO group size $G$ (Eq. 4). To ensure robust conclusions, each configuration is repeated three times for both base and instruct models.
@@ 3.1 Compute-Optimal Scaling
Observation 1 Within the model parameter range of 0.5B to 14B, RL post-training under a fixed computational budget $C$ is compute-optimal when prioritizing the training of larger models rather than extending the training duration of smaller models.
(a) Base Models
(b) Instruct Models
Figure 2: Test loss as a function of training compute ($C$). Larger models consistently achieve lower loss at the same compute budget, showing their higher computational efficiency.
In the compute-constrained setting (Eq. 1), we train 0.5B–14B models and measure test loss as a function of cumulative FLOPs $C$. As shown in Figure 2, larger models consistently outperform smaller ones under the same compute budget for both base and instruct variants. The loss–compute relationship follows a log-linear trend, which can be modeled by a power law:
$\log(L(N,C))=-k_{C}(N)\cdot\log(C)+E_{C}(N).$ (6)
Table 1 shows $k_{C}(N)$ increasing with model size $N$, indicating that larger models not only start from lower loss but also yield greater performance gains per FLOP. This contrasts with pre-training scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), which observe a compute-efficient frontier of the trade-off between model size and computation. In our experiments, no such trade-off is observed, which we believe is due to pre-training allocating substantially greater computational resources and stronger capabilities to larger models that subsequently dominate the post-training dynamics.
This result observed from tested models (0.5B-14B) suggests that, in RL post-training, it is generally preferable to allocate resources toward training larger models. Nevertheless, we hypothesize that this trend may not hold indefinitely, as increasingly larger models start from lower initial loss with making more challenging loss reduction.
Table 1: Fitted slopes $k_{C}$ for different model sizes.
Model 0.5B 1.5B 3B 7B 14B
Base 0.0075 0.0338 0.0168 0.0370 0.0680
Instruct 0.0055 0.0063 0.0153 0.0436 0.0454
@@ 3.2 Data-Optimal Scaling
Observation 2 Within the model range of 0.5B to 14B, for a fixed volume of unique training data $D$, larger models demonstrate superior sample efficiency, consistently achieving lower test loss.
(a) Base Models
(b) Instruct Models
Figure 3: Test loss as a function of data size $D$. Larger models consistently outperform smaller ones given the same number of unique samples.
In the data-constrained setting (Eq. 2), we train models with varying parameter counts on fixed amounts of unique samples. As shown in Figure 3, larger models consistently achieve lower test loss across both base and instruct variants. The loss–data relationship follows a power law:
$\log(L(N,D))=-k_{D}(N)\cdot\log(D)+E_{D}(N)$ (7)
with the sample efficiency coefficient $k_{D}(N)$ increasing with model size (Table 2). It confirms that larger models not only compute more efficiently, but also extract more knowledge from each data point, showing that scaling up model size is the most effective strategy for RL post-training. This observation is similar to pre-training behavior (Kaplan et al., 2020; Hoffmann et al., 2022).
Table 2: Fitted slopes $k_{D}$ for different model sizes.
Model 0.5B 1.5B 3B 7B 14B
Base 0.0070 0.0268 0.0181 0.0409 0.0739
Instruct 0.0054 0.0065 0.0164 0.0469 0.0504

### v2

@@ Abstract
While scaling laws for large language models (LLMs) during pre-training have been extensively studied, their behavior under reinforcement learning (RL) post-training remains largely unexplored. This paper presents a systematic empirical investigation of scaling behaviors in RL-based post-training, with a particular focus on mathematical reasoning. Based on a set of experiments across the full Qwen2.5 dense model series (0.5B to 72B), we characterize how model scale, data volume, and computational budget interact to shape performance. Our analysis leads to four key findings: ❶ Larger models consistently exhibit superior learning efficiency on both compute and data metrics. ❷ The relationship between test loss, compute, and data can be modeled by a predictive power-law which is robust across both base and instruction-tuned models. ❸ Although larger models exhibit higher learning efficiency, the analytical learning efficiency term $k(N)$ in the power-law reveals a latent saturation trend in learning efficiency as model size continues to increase. ❹ In data-constrained regimes, repeated reuse of high-quality data proves highly effective, as final performance is primarily governed by the total number of optimization steps rather than the uniqueness of samples. Collectively, these results provide a principled foundation and practical guidelines for efficiently scaling the reasoning capabilities of LLMs through RL post-training.
@@ 1 Introduction
The rapid progress of large language models (LLMs) has made elucidating their scaling laws a matter of central importance. These laws, which capture the intricate relationships between model architecture, parameter size, computational cost, data availability, and downstream performance (kaplan2020scaling; hoffmann2022training), are invaluable not only because they illuminate the latent factors governing learning dynamics, but also because they provide actionable guidance on how to distribute scarce computational resources most effectively (li2025misfitting). While extensive efforts have clarified scaling behavior, the scaling behavior of reinforcement learning (RL) post-training for LLM reasoning remains underexplored.
@@
During pretraining, kaplan2020scaling show that cross-entropy loss follows smooth power-law scaling in model size, dataset size, and training compute, implying that larger models trained for fewer steps are compute-optimal. hoffmann2022training refine this by showing that, under fixed compute, scaling parameters and tokens proportionally is optimal, since many large models are undertrained.Extending to neural-based RL, hilton2023scaling empirically demonstrates that the intrinsic performance of convolutional neural networks (CNNs) optimized via reinforcement learning also scales like power-law with model capacity and environment interaction.
Recently, RL has become the predominant post-training strategy for enhancing the reasoning abilities of LLMs, particularly in mathematics, a domain that demands long-horizon, compositional reasoning (ferrag2025llm; guo2025deepseekr1; kimi2025scaling; ahn2024large). Given the recent works applying RL for LLM reasoning, it is crucial to understand how to scale RL training effectively for reasoning tasks. In this work, we present a systematic empirical study to characterize these scaling behaviors. We focus on three critical resource regimes: (1) the compute-constrained scenario, where we seek to minimize test loss under a fixed FLOPs budget by determining the optimal model size; (2) the data-constrained scenario, where we identify the model size that yields the lowest test loss given a limited number of unique samples; and (3) the data reuse scenario, where we investigate the trade-off between unique data volume and repeated optimization steps under a fixed computation budget.
Based on our analysis across these regimes, we propose a predictive formulation that characterizes the relationship between test loss $L$, model size NN, and resource budget $X$ (where $X$ denotes either Compute $C$ or Data $D$). We find that the scaling behavior can be effectively modeled by a power law that exhibits a log-linear relationship between test loss and resource budget:
Core Scaling Formula $\log L(N,X)=-k(N)\cdot\log X+E(N).$ (1)
Crucially, our empirical observations suggest that the learning efficiency $k(N)$ does not grow indefinitely with the increasing model scale. Instead, it follows a saturation trend and can be modeled by the formula(Eq. 2) below:
RL Post-Training Learning Efficiency $k(N)=\frac{K_{\text{max}}}{1+\frac{N_{0}}{N}}.$ (2)
This formulation shows that: while larger models consistently exhibit higher learning efficiency (represented by $k(N)$), the marginal gains in efficiency diminish as model size increases, asymptotically approaching a theoretical limit $K_{\text{max}}$.
@@
To empirically validate this law and determine its parameters, we fine-tune 63 LLMs with reinforcement learning on 50k mathematics problems, based on the Qwen2.5 model family (qwen2025qwen25technicalreport). Figure 1 shows that, within the 0.5B-72B range, the loss reduction brought by RL follows an approximately log-linear trend with compute and data size. Importantly, larger models not only have better initial performance but also generally have more efficiency in computation and data utilization during the optimization process. Through deeper analysis, we find that the scaling behavior exhibits notable predictability and further reveal that efficiency gains from model scaling diminish gradually, eventually leading to a saturation effect for larger models.
@@
We further analyze the data-constrained regime, where we demonstrate that data reuse is a highly effective strategy. We validate the generality of our findings through extensive ablation studies on both base and instruct model series. Besides, we have also conducted experiments to study the impact of the rollout number in the GRPO algorithm (shao2024deepseekmathpushinglimitsmathematical), shown in Appendix B.2. These investigations establish fundamental scaling relationships for RL post-training, providing a quantitative foundation and practical guidelines for resource-efficient model refinement.
(a) Relation between test loss and post training compute for base models, fitted on 0.5B-32B and extrapolated on 72B.
(b) Relation between test loss and post training data size for base model, fitted on 0.5B-32B and extrapolated on 72B.
Figure 1: Training Data Fiited on 0.5B-32B with Extrapolation on 72B: In both cases, larger models consistently exhibit higher learning efficiency than smaller models.
@@
  • In our experiment scale, larger models, starting with stronger initial performance, consistently achieve better compute and data efficiency in RL post-training for mathematical reasoning. However, the marginal gains in this efficiency diminish gradually, revealing a saturation trend as model scale increases.
  • The scaling law exhibits predictive capability across both base and instruction-tuned models, allowing us to forecast the training efficiency of larger models and predict the remaining training trajectory from early training data.
@@ 2 Experimental Setup
We use the Qwen2.5 model family (0.5B, 1.5B, 3B, 7B, 14B, 32B and 72B parameters) (qwen2025qwen25technicalreport), which shares the same architecture, so that parameter count is the only variable in our scaling analysis. All experiments are run with the VeRL framework (sheng2024hybridflow), a large-scale RL platform for LLMs ensuring consistency and reproducibility.
Dataset settings.
The training data is the mathematics subset of the guru-RL-92k dataset from the Reasoning360 project (cheng2025revisiting), which is carefully curated through deduplication and difficulty filtering. We further sort the problems by increasing difficulty (decreasing pass rate, evaluated by Qwen2.5-7B-Instruct model) to enable curriculum learning. The evaluation data consists of two parts. To derive scaling laws, we use a held-out set of 500 in-domain math problems sampled from the training distribution. To assess generalization, we evaluate on a broader benchmark suite spanning mathematics (AIME2024 (patel2024aimeaioptimizationmultiple), AMC2023 (knoveleng_amc23_dataset), GSM8K (cobbe2021trainingverifierssolvemath), MATH500 (lightman2023letsverifystepstep)), code (HumanEval (chen2021evaluatinglargelanguagemodels)), logic (Zebra Puzzle(Lin_ZeroEval_A_Unified_2024)), and science (SuperGPQA (pteam2025supergpqascalingllmevaluation)). More details about dataset settings can be found in Appendix A.1.
Prompt Setting.
To ensure stable behavior during RL training and evaluation, we use structured prompts tailored to each domain. For example, all mathematics problems are prepended with the Chain-of-Thought prompt (wei2023chainofthoughtpromptingelicitsreasoning): “You are a knowledgeable math assistant. Answer the following questions and think step by step”. More prompt templates for all related domains could be found in Appendix A.3.
@@
$\mathcal{L}_{\text{GRPO}}=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\lvert o_{i}\rvert}\sum_{t=1}^{\lvert o_{i}\rvert}\Bigl\{\min\Bigl[\rho(\theta)\,\hat{A}_{i,t},\;\operatorname{clip}\!\Bigl(\rho(\theta),1-\varepsilon,\,1+\varepsilon\Bigr)\hat{A}_{i,t}\Bigr]-\beta\,\mathrm{D}_{\mathrm{KL}}\Bigr\},$ (3)
@@
Our primary evaluation metric is the test loss ($L$), a proxy for reward-based performance in the RL setting. Formally, $L=1-(R/R_{\text{max}})$, where $R$ is the number of correct solutions and $R_{\text{max}}$ the total. We adopt the term "test loss" for consistency with foundational neural scaling law literature (kaplan2020scaling). Notably, maximizing reward in RL training is equivalent to minimizing $L$.
Fitting and Prediction Protocols.
To systematically evaluate the robustness and predictive capability of our derived scaling laws, we employ two distinct fitting protocols throughout our analysis:
  • Inter-model Extrapolation: We fit the scaling law parameters using data from smaller models (0.5B to 32B) to calculate the learning efficiency and predict the performance of the larger model (72B).
  • Intra-model Extrapolation: We fit the scaling law using only the early training steps of a specific model to forecast its loss trajectory for the remainder of the training process.
@@ 3 Empirical Results and Scaling Laws
This section presents a comprehensive empirical investigation into the scaling behavior of RL for post-training LLMs. We first examine scaling behaviors under compute and data constraints, then analyze independent scaling dimensions, data reuse strategies, and finally evaluate generalization performance together. To ensure robust conclusions, each configuration is repeated three times for both base and instruct models ranging from 0.5B to 72B.Their statistical uncertainty analysis, including Average Standard Deviation and Standard Error of the Mean (SEM), are provided in Appendix C.3.
@@ 3.1 Compute-Optimal Scaling
To characterize the scaling behavior under computational limits, we first formalize the Compute-Constrained Scenario. Given a fixed computational budget $C$, we seek to identify the optimal model size NN (and the corresponding data allocation $D$) that minimizes the final test loss. This can be expressed as the following constrained optimization problem:
Observation 1 Within the 0.5B–32B parameter range, RL post-training achieves lower test loss under a fixed budget $C$ by prioritizing larger models over extended training duration. However, between 32B and 72B, learning efficiency saturation reduces scaling’s marginal gains, introducing a trade-off between model scale and training steps.
(a) Test loss vs training compute with extrapolation on 72B for base model
(b) Test loss vs training compute with extrapolation on 72B for instruct model
(a) Test loss vs training compute with extrapolation for remainder of each training process on base model
(b) Test loss vs training compute with extrapolation for remainder of each training process on instruct model
Figure 3: Compute Scaling and Predictive Capability from 0.5B-72B for Base and Instruct Models
In the compute-constrained setting, we train 0.5B–72B models and measure test loss as a function of cumulative FLOPs $C$. As shown in Figure 3, larger models consistently outperform smaller ones under the same compute budget for both base and instruct variants. These plots include both Inter-model Extrapolation (fitted on 0.5B-32B and extrapolated on 72B) and Intra-model Prediction (predicting the remainder of training from initial steps) to demonstrate the predictive power of our derived scaling law. The loss–compute relationship follows a log-linear trend, which can be modeled by a power law:
$\log(L(N,C))=-k_{C}(N)\cdot\log(C)+E_{C}(N),\ \text{where}\ \ k_{C}(N)=\left(\frac{K_{Cmax}}{1+\frac{N_{C}}{N}}\right)$ (6)
To demonstrate the predictive capability of the proposed formula Eq 6, we evaluate it in two distinct extrapolation settings:
  1. 1.
    Inter-model Extrapolation: We fit the law’s parameters on smaller models (0.5B–32B) to calculate the learning efficiency ($k_{C}(N)$) of 72B model. As shown in Figure 2(a) and 2(b), the predicted efficiency aligns closely with the actual 72B performance.
  2. 2.
    Intra-model Prediction: We fit the law using only early training steps to forecast the remaining trajectory for a specific model, shown in Figure 3(a) and 3(b).
We further analyze learning efficiency term $k_{C}(N)$ in Eq. 6. As Figure 4 shows, $k_{C}(N)$ grows with model size NN, meaning larger models consistently have higher learning efficiency. However, the efficiency gain from model scale is not uniformly linear. Beyond 32B, the increase in $k_{C}(N)$ diminishes, leading to efficiency saturation.
Figure 4: Fitted learning efficiency $k_{C}(N)$ for Base(Left) and Instruct(Right) models: Larger models consistently exhibit higher learning efficiency with efficiency gains begin to diminish after 32B model.
This saturation is manifested as a distinct performance crossover in Figure 3: In contrast to the immediate dominance of larger models in smaller parameter regimes, the 32B model outperforms the 72B counterpart initially under equivalent compute budgets, as the smaller model size inherently enables more training steps. We believe this observation reveals a latent trade-off between model scale and training steps in compute-constrained scenarios.
@@ 3.2 Data-Optimal Scaling
In many practical applications, the bottleneck lies not in compute but in the availability of high-quality reasoning data. We define the Data-Constrained Scenario as determining the model size NN that yields the lowest test loss given a limited amount of unique training data $D$:
Observation 2 Within the model range of 0.5B to 72B, for a fixed volume of unique training data $D$, larger models demonstrate superior sample efficiency, consistently achieving lower test loss.
(a) Test loss vs data size with extrapolation on 72B for base model
(b) Test loss vs data size with extrapolation on 72B for instruct model
(a) Test loss vs data size with extrapolation for remainder of each training process on base model
(b) Test loss vs data size with extrapolation for remainder of each training process on instruct model
Figure 6: Data Scaling and Predictive Capability from 0.5B-72B for Base and Instruct Models
In the data-constrained setting, we train models with varying parameter counts NN on fixed amounts of unique samples $D$. As shown in Figure 6, larger models consistently achieve lower test loss and higher sample efficiency across both base and instruct variants. The loss–data relationship also follows a log-linear trend and can be modeled by a similar formula addressed in compute scenario:
$\log(L(N,D))=-k_{D}(N)\cdot\log(D)+E_{D}(N),\ \text{where}\ \ k_{D}(N)=\left(\frac{K_{Dmax}}{1+\frac{N_{D}}{N}}\right)$ (8)
Mirroring the analysis in Section 3.1, we evaluate the extrapolative capability of our data scaling law (Eq. 8) in two settings:
  1. 1.
    Inter-model Extrapolation: By fitting parameters on smaller models (0.5B–32B), we accurately predict the data efficiency ($k_{D}(N)$) on 72B model, as illustrated in Figure 5(a) and 5(b).
  2. 2.
    Intra-model Prediction: We forecast the loss trajectory for the remainder of the training process using only early-stage data, shown in Figure 6(a) and 6(b).
We adopt the same analytic form for the data efficiency coefficient $k_{D}(N)$ as we did for compute. As illustrated in Figure 7, $k_{D}(N)$ follows a saturation curve identical to $k_{C}(N)$: while larger models excel at extracting knowledge from each data point, the efficiency gains diminish at scales beyond 32B. The unified functional form across both compute and data domains reflects the theoretical consistency of our scaling law.
Figure 7: Fitted learning efficiency $k_{D}(N)$ for Base(Left) and Instruct(Right) models: $k_{D}(N)$ exhibit nearly identical growth trends with $k_{C}(N)$, with efficiency gains also starting to diminish beyond the 32B model.